OpenStudy (anonymous):

y- intercept of the graph 8x-5y=40

TheSmartOne (thesmartone):

Solve it for y and it should be in the format of y=mx+b where b is the y-intercept.

TheSmartOne (thesmartone):

-5y=-8x+40 y=8/5x-8 y=mx+b So the y-intercept is -8